Still Growing Corp. is a mature firm, growing at a constant rate of 3% per year. Its net income last year was $10 million and it has 2.5 million shares outstanding. Still Growing Corp. has no debt and all income is paid out as dividends to its shareholders. Assuming a 15% cost of capital, what should be the current stock price.
Answer #1

P0 = D0*(1+g) / (k-g)

All income is paid out as dividends, it means the dividend for the last year was $4 per share [ i.e. $10 million / 2.5 million ]

D0 = Dividend for last year = $4

g = growth rate = 3%

k = Cost of equity = 15%

P0 = Current price of the stock

Now let put these figures in the above formula

P0 = D0*(1+g) / (k-g) = 4*(1+3%) / (15%-3%) 34.33